A cylinder rotating at an angular speed of 50rev/sec is brought in contact with an identical stationary cylinder. Because of the kinetic friction, torques act on the two cylinders, acceleration the stationary one and decelerating the moving one. If the common magnitude of the acceleration and deceleration be one revolution per second squared, how long will it take before the two cylinders have equal angular speed?

Grade:11

1 Answers

Arun
5 years ago
Assume , Time = t second.
Init. velocity  = 50 revolution / second.
Fin. velocity = 
In case of first cyl.
ω' = 50 - 1 × t
ω' = 50 - t
In case of second cyl.
Initial speed = 0 and final speed = ω' 
put the value of ω'
50 - t = t × 1
t = 25 second.